OpenStudy (anonymous):

Determine the slope from the table below. x 1, 2, 3, 4, 5 y 1, -8, -17, -26, -35 9 -9 1/9 -8

OpenStudy (anonymous):

\[m = (y_2 - y_1)/(x_2-x_1)\] so you could go with \[(-35-1)/(5-1) = -36/4 = -9\]